OpenStudy (anonymous):

Point P is located at (−2, 7) and point R is located at (1, 0). Find the y value for the point Q that is located two over three the distance from point P to point R.4.9 4.7 (i got this) 2.5 2.3 i just need halp :3 to see if im right...(i kinda guessed) o3o

OpenStudy (anonymous):

@Juakikun @UmarsBack could u two halp meh?

OpenStudy (umarsback):

Lmfaooo i suck at math!!

OpenStudy (anonymous):

lmao XD ik its not my best subject... lmao

OpenStudy (juakikun):

I'm thinking right now if I could remember two years ago of math Lol..

OpenStudy (anonymous):

XD thanks anyway ^-^

OpenStudy (juakikun):

idk if its 2.5 but hell my memory is poor on geo. But honestly its your choice of answer.

OpenStudy (juakikun):

@hartnn can you help her :P?

OpenStudy (anonymous):

should i start off by graphing???? XoX

OpenStudy (anonymous):

i graphed it lol

OpenStudy (anonymous):

@marrero

OpenStudy (anonymous):

....XD

OpenStudy (anonymous):

are you guys all as lost as i am?

hartnn (hartnn):

its much easier after you already graphed it now see the difference in y values -2 and 1

OpenStudy (anonymous):

alright :D yay

hartnn (hartnn):

\(1-(-2)=3\) so the difference is 3 and you need the point at 2.3rd of this distance

hartnn (hartnn):

well, 2/3 of 3 is just 2 so you neead a point at a distance 2 units from A (in your graph)

OpenStudy (anonymous):

ok two units up or down? XD

OpenStudy (anonymous):

i think its 2.5 but im noot sure ight

OpenStudy (anonymous):

alright thanks anyway lol

OpenStudy (anonymous):

hehe no prob

OpenStudy (anonymous):

@hartnn what do u think?

hartnn (hartnn):

instead of going through all steps, lets directly use the formula \(\huge y= \dfrac{my_1+ny_2}{m+n}\) here m : n = 2:3 and y1 = 7 y2 =0

OpenStudy (anonymous):

so \[y=3(7)+2(0) \over 3+2\] ?

OpenStudy (lieutenantgeneral):

4.7

OpenStudy (gudden):

@gingatheninja<3 re-calculate your answer!